OpenStudy (kambol1898):

Match the functions to their x-intercepts. Tiles f(x) = log x – 1 f(x) = -(log x – 2) f(x) = log(-x – 2) f(x) = -log -(x – 1) Pairs (0, 0) arrowBoth (-3, 0) arrowBoth (10, 0) arrowBoth (100, 0) arrowBoth

OpenStudy (anonymous):

to find x-int , set y= 0 \[\log_{10} x - 1 = 0 \] \[\log_{10} x = 1 \] \[x = 10^1 = 10\]

OpenStudy (kambol1898):

thank you but i got it already

OpenStudy (anonymous):

ok , please close thread if you no longer need help

OpenStudy (kambol1898):

i might need your help more often

OpenStudy (kambol1898):

Which exponential function does not have an x-intercept?

OpenStudy (anonymous):

no problem, you can tag me in a question if you like

OpenStudy (anonymous):

All exponential functions do not have x-int unless they have been shifted down \[y = b^x - k, k > 0\]